# Importance of relativistic pericenter precession in identifying the presence of a third body near eccentric binaries

### v1

@@ II.3 Timing model with 1PN corrections
Our goal is to determine the GW bursts timing under the influence of a third body, including the 1PN corrections to the inner binary dynamics. Since 1PN corrections represent the leading order effect that lead to the precession of pericenter, here we focus on the 1PN corrections. However, for completeness, we provide the 2PN corrections to the timing model in Appendix A.
@@
where the first term is the Newtonian acceleration, the second term is the 1PN acceleration and the third term is the leading order tidal perturbation due to third body [^3] In Eq. (13), we neglect the cross term between 1PN acceleration and the third body. This term vanishes for coplanar orbits [83], as we have considered in our study.. The 1PN acceleration is given by [76, 84]
@@
For a highly eccentric orbit, gravitational radiation is primarily emitted during the pericenter passage, while the emission is significantly suppressed during the remainder of the orbit. Assuming that the orbital elements remain constant throughout the orbit and vary only during the pericenter passage, the change in an orbital elements $\bm{x}=[p,e,\omega]$ over a complete orbit is given by
@@
The timing between consecutive bursts can be calculated following a similar procedure. The leading order radiation reaction effects (2.5PN) are incorporated in the timing model by following the Ref. [85]. The complete timing model including the effects of a perturbing third body, the effects of 1PN precession, and 2.5PN radiation reaction effects is given by the following system of equations:
@@
The functions $A(e,p,\eta,m)$ and $B(e)$ in Eq. (II.3) can be found in Eqs. (63)-(66) of Ref. [85]. Note that the tidal corrections to these equations scales as $(m/p)^{-3}\propto v^{-6}$, that corresponds to a -3PN correction. Also note that Eqs. (19) and  (20) only contain 2.5PN effects and third body effects, as there are no 1PN corrections to $p$ and $e$. Equation (21) includes 1PN effects and third body effects, as there are no radiation-reaction corrections to $\omega$. It is important to emphasise that these equations consists of a system of coupled equations, meaning that modifications to any individual equation will affect the others.
@@ III.1 Impact of 1PN corrections on GW burst timing
Figure 3: The difference in GW bursts timings with 1PN ($t$) and without 1PN ($t_{\rm no,1PN}$) corrections as a function of number of burst. Each dot represents the difference in timing recorded at each pericenter passage. The black and red dots denote the unperturbed and perturbed binary, respectively. The horizontal dotted line shows the typical detection threshold ($0.1$ sec) for Einstein Telescope. The initial binary parameters are $m=30M_{\odot},e_{0}=0.99,p_{0}=200m,\omega_{0}=0,V_{3,0}=\pi/3$ perturbed by a third body of mass $m_{3}=10^{6}m$ located at $R=20m_{3}$. For both the unperturbed binary and perturbed binary, the 1PN corrections significantly impact the GW burst timings.
@@
To quantify the effect of 1PN corrections on GW burst timing, we compare the burst times obtained with 1PN corrections to those computed without them. Figure 3 shows the difference in timings with 1PN corrections and without 1PN corrections as a function of the number of bursts. Each dot represents the difference in timing during each pericenter passage. The black dots represent the unperturbed binary and the red dots denote the perturbed binary. For the unperturbed binary, the difference monotonically increases and becomes as large as $\sim 20$ sec at $30^{\rm th}$ burst, which is much greater than the measurement uncertainty in the GW burst timing of 3G detectors $(\sim 0.1\;\text{sec})$ [75]. The perturbed burst timing is initially increasingly delayed compared to the unperturbed case, before reversing and the timing difference becomes significantly negative. The difference becomes as large as $\sim 160$ sec at the end of $30^{\rm th}$ sec (not shown in the plot). This is because $t_{\rm no,1PN}$ starts to increase at a faster rate compared to $t$. As the number of bursts increases further, the mismatch between two timings continues to grow in opposite directions.
Figure 4: The evolution of semi-latus rectum, eccentricity, and longitude of pericenter of the inner binary. The properties of the system are the same as in Fig. 3. Different colors represent different physical scenarios. The perturbed binary with 1PN precession shows distinctive modulation features in $p$ and $e$. The longitude of pericenter increases rapidly with time when 1PN precession is included.
In general, the pronounced difference between the perturbed and unperturbed cases arises from the interplay between perturbative corrections due to the third body and the 1PN acceleration terms for the inner binary. Over a pericenter precession timescale, 1PN effects can accumulate and lead to large shifts in the argument of pericenter, eventually reaching values that are formally non-linear. This transition occurs when $\dot{\omega}_{\rm 1PN}\times t_{\rm obs}\approx 1$ ($\dot{\omega}_{\rm 1PN}$ is the 1PN correction to $\dot{\omega}$ and $t_{\rm obs}$ is the total observation time), corresponding to a significant fraction of a full precession cycle. In this regime, 1PN corrections can exert a disproportionate influence on the GW burst timing, as discussed in [83]. Figure 3 clearly shows this transition: at early times (e.g., before the $\sim 10^{\rm th}$ burst), the system remains in the linear regime, where the 1PN-induced shifts in $\omega$ are relatively small. Beyond this point, the evolution enters the non-linear regime, where cumulative precession effects substantially impact the timing behavior. Therefore, the 1PN corrections significantly impact the burst timing and are crucial for accurately predicting the timings of both perturbed and unperturbed binaries.
@@ III.2 Modulations in the binary parameters
In Figure 4, we plot the binary’s semilatus rectum, eccentricity and longitude of pericenter as a function of time. We choose the same system parameters as in Fig. 3. Different colors represent different physical scenarios. For an unperturbed binary, the evolution of $p$ and $e$ is almost identical. The cyan and the gray dots are on top of each other. With 1PN precession, the longitude of pericenter increases significantly with time for both the perturbed and unperturbed binary. For the perturbed binary, the evolution of $p$ and $e$ is significantly different with and without 1PN precession. In the absence of 1PN precession, $p$ and $e$ do not show any oscillations. However, when 1PN precession is included, it gives rise to periodic oscillations in $p$ and $e$. Notice that in the absence of precession, $p$ increases and $e$ decreases rapidly. When precession is included, both $p$ and $e$ vary within small range. It is important to emphasise that the modulations are only present when the binary is perturbed and 1PN precession is included. Therefore, these modulations in the orbital parameters represent the unique signatures for the presence of a third body near the binary. It is important to note that the magnitude and shape of these modulations depend on the properties of the triple system.
@@
Figure 5: Left: Fractional errors on $m_{3}/R^{3}$ due to neglecting 1PN corrections as a function of $m_{3}/R^{3}$ and for different values of $\omega_{0}$. Other binary parameters are the same as Fig. 3. Right: The errors as a function of $e_{0}$ for different values of $p_{0}$ and $\omega_{0}=0$. The unmodeled 1PN corrections in the timing model leads to significantly large errors in tertiary’s mass and distance.
III.3 Bias in tertiary’s mass and distance due to neglecting 1PN effects
@@
We quantify the systematic errors in third body’s mass and distance due to neglecting 1PN corrections in the timing model. We do so in a simplified way by fitting the model without 1PN corrections to the model with 1PN corrections. In particular, we minimize the square of the difference between the two models summed over the first $40$ GW bursts and varying the combination of $m_{3}/R^{3}$. Since $m_{3}$ and $R$ are highly degenerate parameters, we obtain errors in $m_{3}/R^{3}$. The errors highlight a region of parameter space in which the 1PN corrections are most important.
@@
Figure 5 shows the fractional systematic errors $\Delta(m_{3}/R^{3})/(m_{3}/R^{3})$, where $\Delta(m_{3}/R^{3})$ is the difference between the true and best-fit value of $m_{3}/R^{3}$. Different colors represent various initial values of $\omega_{0}$. We fix the initial values of binary’s parameter to $p_{0}=200m$, $e_{0}=0.99$ and $V_{3,0}=\pi/3$. The errors strongly depend on the initial values of $\omega_{0}$ and can be significantly large, varying from $\mathcal{O}(10^{-1})$ to $\mathcal{O}(10)$. Note that when $\omega_{0}=V_{3,0}=\pi/3$, the errors are $\sim 10$ times larger. This is because when there is no precession, the binary’s pericenter is always aligned with the third object at each orbit, overestimating the influence of the third object. When 1PN precession is included, the pericenter keeps shifting away from the third object, reducing the influence of the third body. Therefore, the errors are largest when $\omega_{0}=V_{3,0}$.
@@
In Fig. 5, we plot the errors in $m_{3}/R^{3}$ as a function of $e_{0}$ for different values of $p_{0}$. We fix $\omega_{0}=0$ in this plot. The errors decrease as the value of $e_{0}$ increases. This is because for fixed values of $p_{0}$, increasing the value of $e_{0}$ makes the binary wider for which the tidal effects of the third body dominate over the 1PN precession. Therefore, errors due to precession decrease as $e_{0}$ increases. Further, when the value of $p_{0}$ decreases for a fixed value of $e_{0}$, the binary becomes more relativistic and 1PN corrections become more important, therefore errors increase significantly. For $e_{0}=0.99$, the errors are $\sim 0.1\mbox{--}10^{2}$. These errors are significantly larger for $e_{0}=0.9$ and range from $\sim 10^{3}\mbox{--}10^{6}$. In conclusion, neglecting 1PN precession leads to significant errors on the mass and distance of the third body. Therefore, it is crucial to incorporate the 1PN corrections in the timing model.
@@ IV Detectability of the perturbing third body
In this section, we discuss the detectability of the perturbing object. To quantify whether a binary is detectable or not, we choose the same detection criterion as in [75] i.e., we compare the difference in the perturbed and unperturbed burst time $(\Delta t)$ with the approximate measurement uncertainity on the burst time $(\sigma_{\rm tb})$. If $\Delta t>\sigma_{\rm tb}$, the tertiary effects are detectable. We choose $\sigma_{\rm tb}\approx 0.1$ which is the mean average standard deviation in burst time posterior samples for Einstein Telescope [75]. Note that $\sigma_{\rm tb}$ depends on the properties of the triple system and the given detector.
Figure 6: The absolute difference between the timings of perturbed and unperturbed binary $|\Delta t|$ as a function of the number of GW burst. The solid lines show the value of $\Delta t$ for a binary with 1PN precession and dashed lines represents the same binary without 1PN precession. The horizontal dotted line indicates the detection threshold. The timings of the first burst are matched so that $\Delta t$ increases with time.We choose an equal mass binary with $m=30M_{\odot},e_{0}=0.99,p_{0}=100m$ and the tertiary of mass $m_{3}=10^{6}m$ at distance of $R=50m_{3}$ and $V_{3,0}=\pi/3$. In the absence of 1PN precession, $\Delta t$ values show significant variation with $\omega_{0}$ and overestimates the phase shift between perturbed and unperturbed binary. The yellow solid line is behind the blue line.
@@ IV.1 The role of 1PN precession on GW burst detectability
Figure 6 shows the absolute difference in the timing $|\Delta t|$ of the same GW burst from the perturbed and the unperturbed binary as a function of the number of GW bursts. The solid lines represent timing with 1PN precession and the dashed lines indicate timing without precession. Different colors represent the different values of $\omega_{0}$. The horizontal dotted line shows $\Delta t=0.1$ as the detection criterion. If $\Delta t>0.1$, the effect of tertiary can be detectable. We choose an equal mass binary with $m=30M_{\odot}$, $e_{0}=0.99$, $p_{0}=100m$, $V_{3,0}=\pi/3$, $m_{3}=10^{6}m$, and $R=50m_{3}$. When precession is not included, the value of $\Delta t$ varies significantly and strongly depends on the initial value of $\omega_{0}$. For some values of $\omega_{0}$, the $\Delta t$ values fall significantly below the detection threshold. However, when the precession is taken into account, the values of $\Delta t$ are very close to each other for different values of $\omega_{0}$. In other words, when precession is taken into account, the dependence of burst timings is weaker on $\omega_{0}$ values. This is because the pericenter precession averages out any dependence on the initial phase of the binary. Note that for the cases when $\Delta t$ is above the detection threshold, unmodeled precession overestimates the value of $\Delta t$ as found in 3-body scattering simulations [70].
Figure 7: The $\Delta t$ as a function of the number of GW bursts. Different colors represent the different values of $R$. The solid lines indicate $\Delta t$ values for $p_{0}=200m$ and dotted lines are for $p_{0}=100m$. Other parameters are the same as Fig. 6.

### v2

@@ II.3 Timing model with 1PN corrections
Our goal is to determine the GW burst timing under the influence of a third body, including the 1PN corrections to the inner binary dynamics. Since 1PN corrections represent the leading order effects that lead to the precession of the pericenter, here we focus on the 1PN corrections. However, for completeness, we provide the 2PN corrections to the timing model in the Appendix A.
@@
where the first term is the Newtonian acceleration, the second term is the 1PN acceleration, and the third term is the leading order tidal perturbation due to the third body. [^4] In Eq. (13), we neglect the cross term between 1PN acceleration and the third body. This term vanishes for coplanar orbits [Will:2013cza], as we have considered in our study. The 1PN acceleration is given by [Poisson_Will_2014, PhysRevD.42.1123]
@@
For a highly eccentric orbit, gravitational radiation is primarily emitted during the pericenter passage, while the emission is significantly suppressed during the remainder of the orbit. Assuming that the orbital elements remain constant throughout the orbit and vary only during the pericenter passage, the secular change in orbital elements $\bm{x}=[p,e,\omega]$ over a complete orbit is given by [Poisson_Will_2014]
@@
The timing between consecutive bursts can be calculated following a similar procedure. The leading order radiation reaction effects (2.5PN) are incorporated in the timing model by following Ref. [Arredondo:2021rdt]. The complete timing model including the effects of a perturbing third body, the effects of 1PN precession, and 2.5PN radiation reaction effects is given by the following system of equations:
@@
The functions $A(e,p,\eta,m)$ and $B(e)$ in Eq. (II.3) can be found in Eqs. (63)-(66) of Ref. [Arredondo:2021rdt]. Note that the tidal correction to these equations scales as $(m/p)^{-3}\propto v^{-6}$, that corresponds to a -3PN correction. Also note that Eqs. (19) and  (20) only contain 2.5PN effects and third body effects, as there are no 1PN corrections to $p$ and $e$. Equation (21) includes 1PN effects and third body effects, as there are no radiation-reaction corrections to $\omega$. It is important to emphasize that these equations consists of a system of coupled equations, meaning that modifications to any individual equation will affect the others.
@@ III.1 Impact of 1PN corrections on GW burst timing
Figure 3: Difference in GW burst timings with 1PN ($t$) and without 1PN ($t_{\rm no,1PN}$) corrections as a function of number of bursts. Each dot represents the difference in timing recorded at each pericenter passage. The black and red dots denote the unperturbed and perturbed binary, respectively. The horizontal dotted line shows the typical detection threshold ($0.1$ sec) for the Einstein Telescope. The initial binary parameters are $m=30M_{\odot},e_{0}=0.99,p_{0}=200m,\omega_{0}=0,V_{3,0}=\pi/3$ perturbed by a third body of mass $m_{3}=10^{6}m$ located at $R=20m_{3}$. For both the unperturbed binary and perturbed binary, the 1PN corrections significantly impact the GW burst timings.
@@
To quantify the effect of 1PN corrections on GW burst timing, we compare the burst times obtained with 1PN corrections to those computed without them. Figure 3 shows the difference in timings with 1PN corrections and without 1PN corrections as a function of the number of bursts. Each dot represents the difference in timing during each pericenter passage. The black dots represent the unperturbed binary, and the red dots denote the perturbed binary. For the unperturbed binary, the difference monotonically increases and becomes as large as $\sim 20$ sec at $30^{\rm th}$ burst, which is much greater than the measurement uncertainty in the GW burst timing of 3G detectors $(\sim 0.1\;\text{sec})$ [Isobel]. The perturbed burst timing is initially increasingly delayed compared to the unperturbed case, before reversing and the timing difference becomes significantly negative. The difference becomes as large as $\sim 160$ sec at the end of the $30^{\rm th}$ sec (not shown in the plot). This is because $t_{\rm no,1PN}$ starts to increase at a faster rate compared to $t$. As the number of bursts increases further, the mismatch between two timings continues to grow in opposite directions.
Figure 4: Evolution of the semi-latus rectum, eccentricity, and longitude of pericenter of the inner binary. The properties of the system are the same as those in Fig. 3. Different colors represent different physical scenarios. The perturbed binary with 1PN precession shows distinctive modulation features in $p$ and $e$. The longitude of the pericenter increases rapidly with time when 1PN precession is included.
In general, the pronounced difference between the perturbed and unperturbed cases arises from the interplay between perturbative corrections due to the third body and the 1PN acceleration terms for the inner binary. Over a pericenter precession timescale, 1PN effects can accumulate and lead to large shifts in the argument of the pericenter, eventually reaching values that are formally non-linear. This transition occurs when $\dot{\omega}_{\rm 1PN}\times t_{\rm obs}\approx 1$ ($\dot{\omega}_{\rm 1PN}$ is the 1PN correction to $\dot{\omega}$ and $t_{\rm obs}$ is the total observation time), corresponding to a significant fraction of a full precession cycle. In this regime, 1PN corrections can exert a disproportionate influence on the GW burst timing, as discussed in [Will:2013cza]. Figure 3 clearly shows this transition: at early times (e.g., before the $\sim 10^{\rm th}$ burst), the system remains in the linear regime, where the 1PN-induced shifts in $\omega$ are relatively small. Beyond this point, the evolution enters the non-linear regime, where cumulative precession effects substantially impact the timing behavior. Therefore, the 1PN corrections significantly impact the burst timing and are crucial for accurately predicting the timings of both perturbed and unperturbed binaries.
@@ III.2 Modulations in the binary parameters
In Figure 4, we plot the binary’s semilatus rectum, eccentricity, and longitude of the pericenter as a function of time. We choose the same system parameters as those in Fig. 3. Different colors represent different physical scenarios. For an unperturbed binary, the evolution of $p$ and $e$ is almost identical. The cyan and the gray dots are on top of each other. With 1PN precession, the longitude of the pericenter increases significantly with time for both the perturbed and unperturbed binary. For the perturbed binary, the evolution of $p$ and $e$ is significantly different with and without 1PN precession. In the absence of 1PN precession, $p$ and $e$ do not show any oscillations. However, when 1PN precession is included, it gives rise to periodic oscillations in $p$ and $e$. Notice that, in the absence of precession, $p$ increases and $e$ decreases rapidly. When precession is included, both $p$ and $e$ vary within a small range. It is important to emphasize that the modulations are only present when the binary is perturbed and 1PN precession is included. Therefore, these modulations in the orbital parameters represent the unique signatures for the presence of a third body near the binary. It is important to note that the magnitude and shape of these modulations depend on the properties of the triple system.
@@
Figure 5: Left: fractional errors on $m_{3}/R^{3}$ due to neglecting 1PN corrections as a function of $m_{3}/R^{3}$ and for different values of $\omega_{0}$. Other binary parameters are the same as those in Fig. 3. Right: errors as a function of $e_{0}$ for different values of $p_{0}$ and $\omega_{0}=0$. The unmodeled 1PN corrections in the timing model lead to significantly large errors in the tertiary mass and distance.
III.3 Bias in the tertiary’s mass and distance due to neglecting 1PN effects
@@
We quantify the systematic errors in the mass and distance of the third body due to neglecting 1PN corrections in the timing model. We do so in a simplified way by fitting the model without 1PN corrections to the model with 1PN corrections. In particular, we minimize the square of the difference between the two models summed over the first $40$ GW bursts and vary the combination of $m_{3}/R^{3}$. Since $m_{3}$ and $R$ are highly degenerate parameters, we obtain errors in $m_{3}/R^{3}$. The errors highlight a region of parameter space in which the 1PN corrections are most important.
@@
Figure 5 shows the fractional systematic errors $\Delta(m_{3}/R^{3})/(m_{3}/R^{3})$, where $\Delta(m_{3}/R^{3})$ is the difference between the true and best-fit value of $m_{3}/R^{3}$. Different colors represent various initial values of $\omega_{0}$. We fix the initial values of the binary parameter to $p_{0}=200m$, $e_{0}=0.99$ and $V_{3,0}=\pi/3$. The errors strongly depend on the initial values of $\omega_{0}$ and can be significantly large, varying from $\mathcal{O}(10^{-1})$ to $\mathcal{O}(10)$. Note that when $\omega_{0}=V_{3,0}=\pi/3$, the errors are $\sim 10$ times larger. This is because when there is no precession, the binary’s pericenter is always aligned with the third object at each orbit, overestimating the influence of the third object. When 1PN precession is included, the pericenter keeps shifting away from the third object, reducing the influence of the third body. Therefore, the errors are largest when $\omega_{0}=V_{3,0}$.
@@
In Fig. 5, we plot the errors in $m_{3}/R^{3}$ as a function of $e_{0}$ for different values of $p_{0}$. We fix $\omega_{0}=0$ in this plot. The errors decrease as the value of $e_{0}$ increases. This is because, for fixed values of $p_{0}$, increasing the value of $e_{0}$ makes the binary wider for which the tidal effects of the third body dominate over the 1PN precession. Therefore, errors due to precession decrease as $e_{0}$ increases. Further, when the value of $p_{0}$ decreases for a fixed value of $e_{0}$, the binary becomes more relativistic and 1PN corrections become more important, and therefore, errors increase significantly. For $e_{0}=0.99$, the errors are $\sim 0.1\mbox{--}10^{2}$. These errors are significantly larger for $e_{0}=0.9$ and range from $\sim 10^{3}$ to $10^{6}$. In conclusion, neglecting 1PN precession leads to significant errors in the mass and distance of the third body. Therefore, it is crucial to incorporate the 1PN corrections in the timing model.
@@ IV Detectability of the perturbing third body
In this section, we discuss the detectability of the perturbing object. To quantify whether a binary is detectable, we choose the same detection criterion as in [Isobel]; i.e., we compare the difference in the perturbed and unperturbed burst times $(\Delta t)$ with the approximate measurement uncertainty on the burst time $(\sigma_{\rm tb})$. If $\Delta t>\sigma_{\rm tb}$, the tertiary effects are detectable. We choose $\sigma_{\rm tb}\approx 0.1$ sec which is the mean average standard deviation in burst time posterior samples for the Einstein Telescope [Isobel]. Note that $\sigma_{\rm tb}$ depends on the properties of the triple system and the given detector.
Figure 6: Absolute difference between the timings of the perturbed and unperturbed binary $|\Delta t|$ as a function of the number of GW bursts. Solid lines show the value of $\Delta t$ for a binary with 1PN precession, and dashed lines represents the same binary without 1PN precession. The horizontal dotted line indicates the detection threshold. The timings of the first burst are matched so that $\Delta t$ increases with time. We choose an equal mass binary with $m=30M_{\odot},e_{0}=0.99,p_{0}=100m$ and a tertiary of mass $m_{3}=10^{6}m$ at distance $R=50m_{3}$ and $V_{3,0}=\pi/3$. In the absence of 1PN precession, $\Delta t$ values show significant variation with $\omega_{0}$ and overestimate the phase shift between the perturbed and unperturbed binary. The yellow solid line is behind the blue line.
@@ IV.1 The role of 1PN precession on GW burst detectability
Figure 6 shows the absolute difference in the timing $|\Delta t|$ of the same GW burst from the perturbed and unperturbed binary as a function of the number of GW bursts. The solid lines represent timing with 1PN precession, and the dashed lines indicate timing without precession. Different colors represent the different values of $\omega_{0}$. The horizontal dotted line shows $\Delta t=0.1$ sec as the detection criterion. If $\Delta t>0.1$ sec, the effect of the tertiary can be detectable. We choose an equal mass binary with $m=30M_{\odot}$, $e_{0}=0.99$, $p_{0}=100m$, $V_{3,0}=\pi/3$, $m_{3}=10^{6}m$, and $R=50m_{3}$. When precession is not included, the value of $\Delta t$ varies significantly and strongly depends on the initial value of $\omega_{0}$. For some values of $\omega_{0}$, the $\Delta t$ values fall significantly below the detection threshold. However, when the precession is taken into account, the values of $\Delta t$ are very close to each other for different values of $\omega_{0}$. In other words, when precession is taken into account, the dependence of burst timings is weaker on $\omega_{0}$ values. This is because the pericenter precession averages out any dependence on the initial phase of the binary. Note that for the cases when $\Delta t$ is above the detection threshold, unmodeled precession overestimates the value of $\Delta t$ as found in 3-body scattering simulations [Samsing:2024syt].
Figure 7: The $\Delta t$ as a function of the number of GW bursts. Different colors represent the different values of $R$. Solid lines indicate $\Delta t$ values for $p_{0}=200m$, and dotted lines are for $p_{0}=100m$. Other parameters are the same as those in Fig. 6.